# Jizhou Liu Booth School of Business

### v1

@@ B.2 Proof of Section 4: Randomization Tests for Weak Null Hypotheses
Following Zhao and Ding (2021), the validity of Procedure 4.1 can be established if we can show the sampling distribution of $T^{WB}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})$ is stochastically dominated by its randomization distribution for almost all sequences of $W$ as $I,J\rightarrow\infty$. To begin with, we provide the following definition of the potential outcomes averaged over the untreated sellers:
$\bar{Y}_{i}^{B}(z)=\frac{1}{J_{0}}\sum_{j=1}^{J}Y_{i,j}(z,0)(1-w_{j}^{S})\text{ for $z\in\{0,1\}$.}$
B.2.1 Sampling Distribution
First, we work on the sampling distribution. Let $\tau$ denote the average buyer spillover effect:
$\tau=\frac{1}{IJ}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(1,0)-\frac{1}{IJ}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(0,0)~{}.$
Let $v_{I}(w^{S})$ denote the finite-population variance of $T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})$ under seller-side assignment $w^{S}$:
$v_{I}(w^{S})=\frac{1}{I_{1}}S_{1}^{2}(w^{S})+\frac{1}{I_{0}}S_{0}^{2}(w^{S})-\frac{1}{I}S_{\tau}^{2}(w^{S})~{},$
where $S_{z}^{2}(w^{S})=(I-1)^{-1}\sum_{i=1}^{I}(\bar{Y}_{i}^{B}(z)-\bar{Y}_{\cdot}^{B}(z))^{2}$ and $S_{\tau}^{2}(w^{S})=(I-1)^{-1}\sum_{i=1}^{I}(\bar{Y}_{i}^{B}(1)-\bar{Y}_{i}^{B}(0)-\tau)^{2}$. Note that, for any fixed $w^{S}$,
$\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})-\tau}{\sqrt{V^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}}\sqrt{\frac{v_{I}(w^{S})+S_{\tau}^{2}(w^{S})/I}{v_{I}(w^{S})}}\leq t\mid w^{S}\right\}-\Phi\left(t\right)\right|$
$=\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})-\tau}{\sqrt{v_{I}(w^{S})}}\frac{\sqrt{v_{I}(w^{S})+S_{\tau}^{2}(w^{S})/I}}{\sqrt{V^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}}\leq t\mid w^{S}\right\}-\Phi(t)\right|$
$=\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})-\tau}{\sqrt{v_{I}(w^{S})}}\leq t\mid w^{S}\right\}-\Phi\left(t\frac{\sqrt{v_{I}(w^{S})+S_{\tau}^{2}(w^{S})/I}}{\sqrt{V^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}}\right)\right|$
$\leq\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})-\tau}{\sqrt{v_{I}(w^{S})}}\leq t\mid w^{S}\right\}-\Phi\left(t\right)\right|+\sup_{t\in\mathbb{R}}\left|\Phi\left(t\frac{\sqrt{v_{I}(w^{S})+S_{\tau}^{2}(w^{S})/I}}{\sqrt{V^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}}\right)-\Phi(t)\right|$ (10)
In Lemma B.1, we show almost sure convergence of $V^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})I$ to $v_{I}(w^{S})I+S_{\tau}^{2}(w^{S})$, which implies that, as $I,J\rightarrow\infty$,
$\sup_{t\in\mathbb{R}}\left|\Phi\left(t\frac{\sqrt{v_{I}(w^{S})+S_{\tau}^{2}(w^{S})/I}}{\sqrt{V^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}}\right)-\Phi\left(t\right)\right|\xrightarrow{a.s.}0$ (11)
Then we define the average treatment effect $\tau(w^{S})$ on the focal units:
$\tau(w^{S})=\frac{1}{IJ_{0}}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(1,0)(1-w^{S}_{j})-\frac{1}{IJ_{0}}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(0,0)(1-w^{S}_{j})=\frac{1}{I}\sum_{i=1}^{I}\bar{Y}_{i}^{B}(1)-\frac{1}{I}\sum_{i=1}^{I}\bar{Y}_{i}^{B}(0)~{}.$
Following Masoero et al. (2024), define $\mathcal{E}_{1,\eta}$ and $\mathcal{E}_{2,\eta}$ as follows:
$\mathcal{E}_{1,\eta}:=\left\{w^{S}:\left|v_{I}(w^{S})-\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}\right|\leq\Delta_{1,\eta}\right\}$
$\mathcal{E}_{2,\eta}:=\left\{w^{S}:|\tau(w^{S})-\tau|\leq\Delta_{2,\eta}\right\}~{},$
@@ B.2.1 Sampling Distribution
$\Delta_{1,\eta}:=CC_{2}^{2}\left[\Delta_{I}\Delta_{J}+\Delta_{J}^{2}\right]\log(1/\eta)$
$\Delta_{2,\eta}:=CC_{2}\Delta_{J}\log(C^{\prime}/\eta)$
$\Delta_{I}:=I_{0}^{-1}+I_{1}^{-1}$
$\Delta_{J}:=J_{0}^{-1}+J_{1}^{-1}~{}.$
Note that
$\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})-\tau}{\sqrt{v_{I}(w^{S})}}\leq t\mid w^{S}\right\}-\Phi\left(t\right)\right|$
$=\sup_{t\in\mathbb{R}}\left|P\left\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\leq t\mid w^{S}\right\}-\Phi\left(\frac{t-\tau}{\sqrt{v_{I}(w^{S})}}\right)\right|$
$\leq\sup_{t\in\mathbb{R}}\left|P\left\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\leq t\mid w^{S}\right\}-\Phi\left(\frac{t-\tau(w^{S})}{\sqrt{v_{I}(w^{S})}}\right)\right|+\sup_{t\in\mathbb{R}}\left|\Phi\left(\frac{t-\tau(w^{S})}{\sqrt{v_{I}(w^{S})}}\right)-\Phi\left(\frac{t-\tau}{\sqrt{v_{I}(w^{S})}}\right)\right|$
$\leq\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})-\tau(w^{S})}{\sqrt{v_{I}(w^{S})}}\leq t\mid w^{S}\right\}-\Phi\left(t\right)\right|+\sup_{t\in\mathbb{R}}\left|\Phi\left(\frac{t-\tau(w^{S})}{\sqrt{v_{I}(w^{S})}}\right)-\Phi\left(\frac{t-\tau}{\sqrt{v_{I}(w^{S})}}\right)\right|~{}.$
By Lemma B.5 of Masoero et al. (2024), given $w^{S}\in\mathcal{E}_{1,\eta}$, we have
$\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})-\tau(w^{S})}{\sqrt{v_{I}(w^{S})}}\leq t\mid w^{S}\right\}-\Phi\left(t\right)\right|\leq\frac{C\left(\Delta_{I}C_{2}+\sqrt{\Delta_{1,\eta}}\right)}{\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}^{\frac{1}{2}}},$
Next, note that
$\sup_{t\in\mathbb{R}}\left|\Phi\left(\frac{t-\tau(w^{S})}{\sqrt{v_{I}(w^{S})}}\right)-\Phi\left(\frac{t-\tau}{\sqrt{v_{I}(w^{S})}}\right)\right|$ $\leq\sup_{t\in\mathbb{R}}\left|\Phi\left(\frac{t-\tau(w^{S})}{\sqrt{v_{I}(w^{S})}}\right)-\Phi\left(\frac{t-\tau(w^{S})}{\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}^{\frac{1}{2}}}\right)\right|$
$\hskip 30.00005pt+\sup_{t\in\mathbb{R}}\left|\Phi\left(\frac{t-\tau}{\sqrt{v_{I}(w^{S})}}\right)-\Phi\left(\frac{t-\tau(w^{S})}{\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}^{\frac{1}{2}}}\right)\right|$
Following Lemma B.17 in Masoero et al. (2024), for any $\eta\leq 1$, we have
$\sup_{t\in\mathbb{R}}\left|\Phi\left(\frac{t-\tau(w^{S})}{\sqrt{v_{I}(w^{S})}}\right)-\Phi\left(\frac{t-\tau(w^{S})}{\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}^{\frac{1}{2}}}\right)\right|$
$\leq\eta+\frac{2|\sqrt{v_{I}(w^{S})}-\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}^{\frac{1}{2}}|\sqrt{\log(e/\eta)}}{\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}^{\frac{1}{2}}}$

### v2

@@ B.2 Proof of Section 4: Randomization Tests for Weak Null Hypotheses
This section provides the proofs of Theorem 4.1 and 4.2 and its supporting lemmas. Throughout this section, we repeatedly use large-sample results for Simple Random Sampling Without Replacement (SRSWOR) to handle the dependencies inherent in finite-population inference.
B.2.1 Proof of Theorem 4.1
Proof.
The asymptotic validity of Procedure 4.1 under $H_{0}^{wb,2}$ and its potential invalidity under $H_{0}^{wb,1}$ follow from the limit results established in Lemma B.3.
First, consider the null hypothesis $H_{0}^{wb,2}$ in (9). Under this hypothesis, $\tau(w^{S})=\tau=0$ by construction, causing the first stochastic term $A_{N}$ to vanish. Lemma B.3 then implies that the sampling distribution of the studentized statistic converges to a normal distribution with variance $\sigma_{B}^{2}\leq 1$:
$T^{WB}\Rightarrow N(0,\sigma_{B}^{2}),\quad\sigma_{B}^{2}\leq 1.$
Concurrently, the conditional randomization distribution used in Procedure 4.1—which is constructed by permuting $w^{B}$ while holding $w^{S}$ fixed—converges in probability to a standard normal distribution, $N(0,1)$, following the general results for studentized statistics under complete randomization (Wu and Ding, 2021; Zhao and Ding, 2021). Because the limit of the randomization distribution has a variance of 1, which is greater than or equal to the sampling variance $\sigma_{B}^{2}$, the randomization distribution stochastically dominates the sampling distribution. This ensures that the test is asymptotically conservative, satisfying $\lim\Pr(\text{Reject})\leq\alpha$.
In contrast, under the global weak null $H_{0}^{wb,1}$ in (8), Lemma B.3 shows that the sampling distribution converges to $N(0,\sigma_{A}^{2}+\sigma_{B}^{2})$. As established in the subsequent remarks, the total variance $\sigma_{A}^{2}+\sigma_{B}^{2}$ can strictly exceed 1 depending on the values of $S_{\mu^{\Delta}}^{2}$ and $\lambda$. In such cases, the sampling distribution is “wider” than the $N(0,1)$ randomization distribution, leading to a systematic over-rejection of the null hypothesis and thus a loss of asymptotic validity.
B.2.2 Finite-Population Central Limit Theorems
We now state the regularity conditions required for our asymptotic theory under two-sided experiments with complete randomization. We consider a sequence of finite populations indexed by $N$, where $I=I_{N}$ and $J=J_{N}$. Let $p_{1}:=I_{1}/I\to\pi_{1}\in(0,1)$, $p_{0}:=I_{0}/I\to\pi_{0}\in(0,1)$, $p_{2}:=J_{1}/J\to\pi_{2}\in(0,1)$, and assume $I\to\infty,J\to\infty,I/J\to\lambda\in(0,\infty)$.
Assumption B.1 (Bounded fourth moments).
The potential outcomes satisfy
$\sup_{N}\frac{1}{IJ}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{ij}(z,0)^{4}<\infty\quad(z\in\{0,1\}),\qquad\sup_{N}\frac{1}{IJ}\sum_{i=1}^{I}\sum_{j=1}^{J}\Delta_{ij}^{4}<\infty,$
where $\Delta_{ij}:=Y_{ij}(1,0)-Y_{ij}(0,0)$.
Assumption B.2 (Moment control for seller-averaging error).
For each $z\in\{0,1\}$, let
$\mu_{i}(z):=\frac{1}{J}\sum_{j=1}^{J}Y_{ij}(z,0),\qquad S_{i,z}^{2}:=\frac{1}{J-1}\sum_{j=1}^{J}\big(Y_{ij}(z,0)-\mu_{i}(z)\big)^{2},$
and assume
$\sup_{N}\frac{1}{I}\sum_{i=1}^{I}S_{i,z}^{2}<\infty\quad(z\in\{0,1\}).$
Assumption B.3 (Buyer-level nondegeneracy).
Define $\delta_{i}:=\mu_{i}(1)-\mu_{i}(0)$ and the finite-population variances
$S_{z}^{2}:=\frac{1}{I-1}\sum_{i=1}^{I}\big(\mu_{i}(z)-\bar{\mu}(z)\big)^{2},\quad z\in\{0,1\},\qquad S_{\delta}^{2}:=\frac{1}{I-1}\sum_{i=1}^{I}(\delta_{i}-\bar{\delta})^{2},$
where $\bar{\mu}(z):=I^{-1}\sum_{i=1}^{I}\mu_{i}(z)$ and $\bar{\delta}:=I^{-1}\sum_{i=1}^{I}\delta_{i}$. Assume $S_{0}^{2},S_{1}^{2}$ converge to finite limits with $\liminf_{N}S_{0}^{2}>0$ and $\liminf_{N}S_{1}^{2}>0$.
Assumption B.4 (Uniform integrability for Lindeberg conditions).
The following uniform integrability conditions hold:
  1. (i)
    For the seller-level treatment effects $\mu_{j}^{\Delta}:=I^{-1}\sum_{i=1}^{I}\Delta_{ij}$,
    $\lim_{M\to\infty}\sup_{N}\frac{1}{J}\sum_{j=1}^{J}(\mu_{j}^{\Delta})^{2}\mathbf{1}\{(\mu_{j}^{\Delta})^{2}>M\}=0.$
  2. (ii)
    For each $z\in\{0,1\}$, conditional on any seller assignment $w^{S}$ with $|S_{0}(w^{S})|=J_{0}\asymp J$,
    $\lim_{M\to\infty}\sup_{N}E\left[\frac{1}{I}\sum_{i=1}^{I}\big(\bar{Y}_{i}^{B}(z;w^{S})\big)^{2}\mathbf{1}\big\{\big(\bar{Y}_{i}^{B}(z;w^{S})\big)^{2}>M\big\}\right]=0.$
Remark B.1.
Assumptions B.1-B.4 are natural conditions required for finite-population central limit theorems. The boundedness conditions in Assumptions B.1, B.2, and B.4 are satisfied if the potential outcomes are bounded, as assumed in Masoero et al. (2026). The nondegeneracy condition in Assumption B.3 is also a standard assumption for the finite-population CLT (Li and Ding, 2017). As we will show later, Assumption B.3 is essential for ensuring the validity of the studentization method based on our proposed two-way variance estimator.
The following lemmas establish the necessary central limit theorems for each side of randomization. These results follow from the general finite-population CLT theory developed by Li and Ding (2017). The key to applying these results is verifying the Lindeberg-type conditions required by Li and Ding (2017).
Lemma B.1 (Seller-side SRSWOR CLT).
Let $\mu_{j}^{\Delta}:=I^{-1}\sum_{i=1}^{I}\Delta_{ij}$ and define the finite-population variance
$S_{\mu^{\Delta}}^{2}:=\frac{1}{J-1}\sum_{j=1}^{J}(\mu_{j}^{\Delta}-\tau)^{2},$
@@ B.2.2 Finite-Population Central Limit Theorems
$\tau:=\frac{1}{IJ}\sum_{i=1}^{I}\sum_{j=1}^{J}\big\{Y_{ij}(1,0)-Y_{ij}(0,0)\big\}.$
Assume $S_{\mu^{\Delta}}^{2}$ converges to a finite positive limit. Under Assumptions B.1B.4 and the null hypothesis $H_{0}^{wb}:\tau=0$, we have
$\sqrt{J}\tau(w^{S})=\sqrt{J}\cdot\frac{1}{J_{0}}\sum_{j:w_{j}^{S}=0}\mu_{j}^{\Delta}\Rightarrow N\!\left(0,\ \frac{\pi_{2}}{1-\pi_{2}}\,S_{\mu^{\Delta}}^{2}\right),$
where $S_{0}(w^{S}):=\{j\in[J]:w_{j}^{S}=0\}$ is an SRSWOR subset of $[J]$ of size $J_{0}$.
Proof.
The seller assignment $w^{S}$ induces an SRSWOR of size $J_{0}$ from the finite population $\{\mu_{j}^{\Delta}\}_{j=1}^{J}$. Under $H_{0}^{wb,1}$, the population mean is $\tau=0$.
Define the SRSWOR sample mean $\tau(w^{S})=J_{0}^{-1}\sum_{j:w_{j}^{S}=0}\mu_{j}^{\Delta}$ and the finite-population variance $S_{\mu^{\Delta}}^{2}=(J-1)^{-1}\sum_{j=1}^{J}(\mu_{j}^{\Delta})^{2}$ (using $\tau=0$).
Step 1: Verify conditions for Li and Ding (2017) Theorem 1.
The finite-population CLT for SRSWOR in Li and Ding (2017) (Theorem 1) requires:
  1. 1.
    Sample size growth: $J_{0}\to\infty$ and $J-J_{0}\to\infty$.
    This is ensured by $J\to\infty$ and $J_{1}/J\to\pi_{2}\in(0,1)$, which implies $J_{0}=J-J_{1}\sim(1-\pi_{2})J\to\infty$.
  2. 2.
    Lindeberg condition:
    $\frac{\max_{j\in[J]}(\mu_{j}^{\Delta}-\tau)^{2}}{J\,S_{\mu^{\Delta}}^{2}}\to 0.$ (14)
    Under $H_{0}^{wb,1}$, $\tau=0$, so this becomes
    $\frac{\max_{j\in[J]}(\mu_{j}^{\Delta})^{2}}{J\,S_{\mu^{\Delta}}^{2}}\to 0.$
Step 2: Derive the Lindeberg condition from Assumption B.4.
We need to verify the Lindeberg condition:
$\frac{\max_{j\in[J]}(\mu_{j}^{\Delta}-\tau)^{2}}{J\,S_{\mu^{\Delta}}^{2}}\to 0.$
Under $H_{0}^{wb,1}$, $\tau=0$, so this becomes
$\frac{\max_{j\in[J]}(\mu_{j}^{\Delta})^{2}}{J\,S_{\mu^{\Delta}}^{2}}\to 0.$
Since $S_{\mu^{\Delta}}^{2}$ converges to a positive limit $\sigma^{2}>0$, we have $S_{\mu^{\Delta}}^{2}\geq\sigma^{2}/2$ for sufficiently large $N$. Therefore, it suffices to show
$\frac{\max_{j\in[J]}(\mu_{j}^{\Delta})^{2}}{J}\to 0.$
For any $\epsilon>0$, we have
$\frac{1}{J}\max_{j\in[J]}(\mu_{j}^{\Delta})^{2}\leq\frac{1}{J}\sum_{j=1}^{J}(\mu_{j}^{\Delta})^{2}\mathbf{1}\{(\mu_{j}^{\Delta})^{2}>\epsilon J\}+\epsilon.$
For the first term, by Assumption B.4(i) with $M=\epsilon J$:
$\sup_{N}\frac{1}{J}\sum_{j=1}^{J}(\mu_{j}^{\Delta})^{2}\mathbf{1}\{(\mu_{j}^{\Delta})^{2}>\epsilon J\}\leq\sup_{N}\frac{1}{J}\sum_{j=1}^{J}(\mu_{j}^{\Delta})^{2}\mathbf{1}\{(\mu_{j}^{\Delta})^{2}>M\}\to 0$
as $M=\epsilon J\to\infty$ (since $J\to\infty$).
Therefore,
$\limsup_{N\to\infty}\frac{1}{J}\max_{j\in[J]}(\mu_{j}^{\Delta})^{2}\leq\epsilon.$
Since $\epsilon>0$ was arbitrary, we conclude
$\frac{\max_{j\in[J]}(\mu_{j}^{\Delta})^{2}}{J}\to 0,$
which establishes the Lindeberg condition (14).
Step 3: Apply Li and Ding (2017) Theorem 1.
With both conditions verified, Li and Ding (2017) Theorem 1 yields
$\frac{\sqrt{J_{0}}\,\tau(w^{S})}{\sqrt{(1-f_{J})S_{\mu^{\Delta}}^{2}}}\Rightarrow N(0,1),$
where $f_{J}:=J_{0}/J$ is the sampling fraction. Since $J_{0}=J-J_{1}$ and $J_{1}/J\to\pi_{2}$, we have $f_{J}\to 1-\pi_{2}$ and thus $1-f_{J}\to\pi_{2}$.
Rearranging gives
$\sqrt{J_{0}}\,\tau(w^{S})\Rightarrow N\!\left(0,\ \pi_{2}S_{\mu^{\Delta}}^{2}\right).$
Multiplying both sides by $\sqrt{J/J_{0}}$ and using $J/J_{0}\to 1/(1-\pi_{2})$ yields
$\sqrt{J}\,\tau(w^{S})\Rightarrow N\!\left(0,\ \frac{\pi_{2}}{1-\pi_{2}}\,S_{\mu^{\Delta}}^{2}\right),$
completing the proof.
For notational brevity, we suppress the superscript “obs” and write
$(w^{B},w^{S})=\mathbf{W}^{obs},\qquad\mathbf{Y}^{obs}=\mathbf{Y}(\mathbf{W}^{obs}),$
and we abbreviate
$T^{B}(w^{B},w^{S})$ $:=T^{B}(\mathbf{W}^{obs}\mid\mathbf{Y}^{obs},\mathcal{C})$
$V^{B}(w^{B},w^{S})$ $:=V^{B}(\mathbf{W}^{obs}\mid\mathbf{Y}^{obs},\mathcal{C})$
$T^{WB}(w^{B},w^{S})$ $:=T^{WB}(\mathbf{W}^{obs}\mid\mathbf{Y}^{obs},\mathcal{C})~.$
Lemma B.2 (Buyer-side CRD CLT conditional on seller assignment).
Conditional on the seller assignment $w^{S}$, define the seller-focal buyer-level potential outcomes
$\bar{Y}_{i}^{B}(z;w^{S}):=\frac{1}{J_{0}}\sum_{j=1}^{J}Y_{ij}(z,0)\,(1-w_{j}^{S}),\qquad\bar{Y}_{\cdot}^{B}(z;w^{S}):=\frac{1}{I}\sum_{i=1}^{I}\bar{Y}_{i}^{B}(z;w^{S}),$
and the seller-focal average treatment effect
$\tau(w^{S}):=\bar{Y}_{\cdot}^{B}(1;w^{S})-\bar{Y}_{\cdot}^{B}(0;w^{S}).$
Let the difference-in-means statistic be
$T^{B}(w^{B},w^{S}):=\frac{1}{I_{1}}\sum_{i=1}^{I}w_{i}^{B}\,\bar{Y}_{i}^{B}(1;w^{S})-\frac{1}{I_{0}}\sum_{i=1}^{I}(1-w_{i}^{B})\,\bar{Y}_{i}^{B}(0;w^{S}).$
Define the conditional finite-population variance
$v_{I}(w^{S}):=\frac{S_{1}^{2}(w^{S})}{I_{1}}+\frac{S_{0}^{2}(w^{S})}{I_{0}}-\frac{S_{\tau}^{2}(w^{S})}{I},$
where
$S_{z}^{2}(w^{S}):=\frac{1}{I-1}\sum_{i=1}^{I}\big(\bar{Y}_{i}^{B}(z;w^{S})-\bar{Y}_{\cdot}^{B}(z;w^{S})\big)^{2},$
$\delta_{i}(w^{S}):=\bar{Y}_{i}^{B}(1;w^{S})-\bar{Y}_{i}^{B}(0;w^{S}),\qquad S_{\tau}^{2}(w^{S}):=\frac{1}{I-1}\sum_{i=1}^{I}\big(\delta_{i}(w^{S})-\tau(w^{S})\big)^{2}.$
Under Assumptions B.1B.4, conditional on $w^{S}$, the centered difference-in-means satisfies
$\frac{T^{B}(w^{B},w^{S})-\tau(w^{S})}{\sqrt{v_{I}(w^{S})}}\Rightarrow N(0,1).$
Proof.
Conditional on $w^{S}$, the buyer assignment $w^{B}$ is a two-arm complete randomization with fixed margins $(I_{1},I_{0})$ applied to the finite population of seller-focal potential outcomes $\{\bar{Y}_{i}^{B}(0;w^{S}),\bar{Y}_{i}^{B}(1;w^{S})\}_{i=1}^{I}$.
Step 1: Verify conditions for Li and Ding (2017).
The finite-population CLT for complete randomization in Li and Ding (2017) requires:
  1. 1.
    Sample size growth: $I_{1}\to\infty$ and $I_{0}\to\infty$.
    This is ensured by $I\to\infty$ and $I_{1}/I\to\pi_{1}\in(0,1)$.
  2. 2.
    Lindeberg condition: For each $z\in\{0,1\}$,
    $\frac{\max_{i\in[I]}\big(\bar{Y}_{i}^{B}(z;w^{S})-\bar{Y}_{\cdot}^{B}(z;w^{S})\big)^{2}}{I\,S_{z}^{2}(w^{S})}\xrightarrow{p}0.$ (15)
  3. 3.
    Nondegeneracy: $S_{z}^{2}(w^{S})$ does not degenerate, which will be established in the proof of the main lemma (Lemma B.3, Step 1(e)) via Assumption B.3.
Step 2: Derive the Lindeberg condition from Assumption B.4.
Fix $z\in\{0,1\}$. We need to verify the Lindeberg condition:
$\frac{\max_{i\in[I]}\big(\bar{Y}_{i}^{B}(z;w^{S})-\bar{Y}_{\cdot}^{B}(z;w^{S})\big)^{2}}{I\,S_{z}^{2}(w^{S})}\xrightarrow{p}0.$ (16)
Since $S_{z}^{2}(w^{S})\xrightarrow{p}S_{z}^{2}>0$ (established in Step 1(e) of the main lemma), it suffices to show
$\frac{\max_{i\in[I]}\big(\bar{Y}_{i}^{B}(z;w^{S})-\bar{Y}_{\cdot}^{B}(z;w^{S})\big)^{2}}{I}\xrightarrow{p}0.$
Using the inequality $\max_{i}(x_{i}-\bar{x})^{2}\leq 2\max_{i}x_{i}^{2}+2\bar{x}^{2}$, we have
$\max_{i\in[I]}\big(\bar{Y}_{i}^{B}(z;w^{S})-\bar{Y}_{\cdot}^{B}(z;w^{S})\big)^{2}\leq 2\max_{i\in[I]}\big(\bar{Y}_{i}^{B}(z;w^{S})\big)^{2}+2\big(\bar{Y}_{\cdot}^{B}(z;w^{S})\big)^{2}.$
For the first term, for any $\epsilon>0$:
$\frac{1}{I}\max_{i\in[I]}\big(\bar{Y}_{i}^{B}(z;w^{S})\big)^{2}\leq\frac{1}{I}\sum_{i=1}^{I}\big(\bar{Y}_{i}^{B}(z;w^{S})\big)^{2}\mathbf{1}\big\{\big(\bar{Y}_{i}^{B}(z;w^{S})\big)^{2}>\epsilon I\big\}+\epsilon.$
By Assumption B.4(ii) with $M=\epsilon I$, the first term vanishes as $N\to\infty$ (since $I\to\infty$). Thus
$\limsup_{N\to\infty}\frac{1}{I}\max_{i\in[I]}\big(\bar{Y}_{i}^{B}(z;w^{S})\big)^{2}\leq\epsilon.$
Since $\epsilon>0$ was arbitrary,
$\frac{1}{I}\max_{i\in[I]}\big(\bar{Y}_{i}^{B}(z;w^{S})\big)^{2}\xrightarrow{p}0.$
For the second term, $\bar{Y}_{\cdot}^{B}(z;w^{S})^{2}=O_{p}(1)$ (shown in Step 1(c) of the main lemma), so $\bar{Y}_{\cdot}^{B}(z;w^{S})^{2}/I\to 0$ in probability.
Combining these results establishes (16).
Step 3: Apply Li and Ding (2017).
With all conditions verified, Li and Ding (2017) yields the stated result. The normalizing variance $v_{I}(w^{S})$ is precisely the Neyman conservative variance formula for completely randomized experiments, which is the form given in Li and Ding (2017) for two-arm CRD.
B.2.3 CLT for Neyman-style Studentized Statistic
We now present the main asymptotic result for the studentized buyer-spillover statistic under two-sided complete randomization.
Lemma B.3 (Two-term CLT for Neyman-studentized buyer-spillover statistic under two-sided complete randomization).
Let there be $I=I_{N}$ buyers and $J=J_{N}$ sellers. For each buyer–seller pair $(i,j)$ and each $(b,s)\in\{0,1\}^{2}$, fix dyad-level potential outcomes $Y_{ij}(b,s)\in\mathbb{R}$.
Two-sided complete randomization with fixed margins. Draw the seller assignment $w^{S}=(w_{1}^{S},\ldots,w_{J}^{S})\in\{0,1\}^{J}$ uniformly over all vectors with exactly $J_{1}$ ones and $J_{0}:=J-J_{1}$ zeros. Independently draw the buyer assignment $w^{B}=(w_{1}^{B},\ldots,w_{I}^{B})\in\{0,1\}^{I}$ uniformly over all vectors with exactly $I_{1}$ ones and $I_{0}:=I-I_{1}$ zeros.
Write $p_{1}:=I_{1}/I\to\pi_{1}\in(0,1)$, $p_{0}:=I_{0}/I\to\pi_{0}\in(0,1)$, $p_{2}:=J_{1}/J\to\pi_{2}\in(0,1)$, and assume
$I\to\infty,\qquad J\to\infty,\qquad\frac{I}{J}\to\lambda\in[0,\infty).$
Neyman variance and studentized statistic. Define the Neyman (conservative) variance estimator
$V^{B}(w^{B},w^{S}):=\frac{s_{1}^{2}(w^{B},w^{S})}{I_{1}}+\frac{s_{0}^{2}(w^{B},w^{S})}{I_{0}},$
where $s_{1}^{2}(w^{B},w^{S})=s_{1}^{2}$ and $s_{0}^{2}(w^{B},w^{S})=s_{0}^{2}$ with $s_{1}^{2},s_{0}^{2}$ defined in Procedure 4.1, as we use notation $(w^{B},w^{S})$ to emphasize the dependence on buyer and seller side randomization.
Define the studentized statistic
$T^{WB}(w^{B},w^{S}):=\frac{T^{B}(w^{B},w^{S})}{\sqrt{V^{B}(w^{B},w^{S})}}.$
Two-term decomposition. Define
$T^{WB}(w^{B},w^{S})=A_{N}+B_{N},\qquad A_{N}:=\frac{\tau(w^{S})}{\sqrt{V^{B}(w^{B},w^{S})}},\quad B_{N}:=\frac{T^{B}(w^{B},w^{S})-\tau(w^{S})}{\sqrt{V^{B}(w^{B},w^{S})}}.$
Asymptotic variances. Define the deterministic limit
$D:=\frac{S_{1}^{2}}{\pi_{1}}+\frac{S_{0}^{2}}{\pi_{0}},$